# UNGT: Ultrasound Nasogastric Tube Dataset for Medical Image Analysis

### v1

@@ 3.1 Data Acquisition and Characteristics
To mitigate the gap from the dataset aspect, we propose a UNGT dataset with 493 US images gathered from 110 patients. Of these, 305 images are captured in the transverse section while 188 images are scanned in the longitudinal section. The average resolution of the scans is approximately 879 $\times$ 583. The UNGT dataset comprises precise annotations for the liver, stomach, tube, and pancreas and we partition it into two components. Specifically, component A encompasses 321 images with the tube collected from 60 patients, and component B consists of 172 images without the tube gathered from 50 patients. We demonstrate several example images within the developed UNGT dataset in fig. 1a. For neatness consideration, images have been resized and cropped as discussed in Section 4.1.
@@
We construct the dataset in two approaches. First, we retrospectively trawl the institution imaging database PACS between 1 January 2020 and 31 December 2022 to retrieve pre-existing US images demonstrating the stomach with or without the feeding tube. Second, we prospectively recruit patients with the pre-existing feeding tube in situ for US scanning after the approval of the ethics review board and the acquisition of written patient consent. The IRB number for our work is 2022/00171. We perform the scans using standard commercial US machines, including General Electric LOGIQ E9 and E10, Siemens ACUSON, and Toshiba APLIO 100 on transverse and longitudinal sections. We set the curvilinear probe frequency to 2–8 MHz based on the habitus and depth of the region of interest for optimal image contrast and resolution. We obtain approximately 2 to 10 static images per patient and capture more frames when the cine-clips are applicable. We review the quality of the US images to ensure adequate clarity, resolution, and field-of-view before selection. A board-certified radiologist with 10 years of experience in US imaging selects the scans, performs cropping to remove confidential information, and carries out the annotation using an open-source tool Labelme [45]. The liver, stomach, tube, and pancreas are annotated, with the tube consistently inside the stomach. Another board-certified radiologist with 5 years of experience in US imaging cross-checks the annotations for accuracy. Annotation discrepancies are resolved by consensus between the two radiologists.
@@
We present the characteristics analysis of our developed UNGT dataset in fig. 1b to fig. 1e. The distribution of image counts per patient reveals that only a limited number of images are typically captured per patient, highlighting the challenges of US image acquisition. The brightness histogram distribution aligns with the typical characteristics of US images, which often exhibit predominantly low average brightness. We exclude the grey level of zero as it can comprise non-scanned regions, such as black triangle borders outside the scan boundary. The primary challenge within the UNGT dataset is the severely unbalanced distribution. Considering the image counts per structure, most images comprise the liver and stomach while only a small portion involve the tube and pancreas. Such an unbalanced constraint becomes more severe for the size across varying structures. The number of pixels labeled as liver and stomach is not at the same magnitude as that of the tube and pancreas. This can be attributed to the limited number of structures and the minor size of such objects. The severely unbalanced distribution at coarse and fine levels demonstrates the difficulty in analyzing the UNGT dataset, especially for the tube and pancreas structures.
@@ 3.2 Model Architecture Design
Figure 2: Architecture of the developed AAMS. It comprises an ADW scheme, an MAA block, the OAA, and the RRE blocks. The MAA block is mainly composed of a local excitation (LEX) block, a global excitation (GEX) block, and a multi-layer perceptron (MLP) block. $\oplus$ and $\times$ denote the exclusive OR and multiplication operations. L, S, T, and P stand for the liver, stomach, tube, and pancreas. HAAP and VAAP represent horizontal and vertical adaptive average pooling layers. DwConv, SpConv, and PwConv present depthwise, stripe, and pointwise convolution layers. BN stands for batch normalization layer. (a) AAMS architecture, (b) ADW scheme, and (c) MAA block.
@@
We advance the DEMS to the multi-category segmentation context using parallel binary segmentation and propose an ADW approach and an MAA block. We present the detailed architecture of the proposed AAMS in fig. 2. It comprises an ADW scheme, an MAA block, the OAA, and the RRE blocks. The ADW algorithm tackles class imbalance by adaptively regulating the loss contributions across various categories during training. The MAA block enhances feature representation by integrating local and global contextual representations. The loss function of AAMS comprises fusion loss $L_{f}$, sensitivity loss $L_{s}$, and unsupervised loss $L_{u}$.
The ADW computes the number of pixels for each category for each batch. Given a batch of encoded mask $M$, the labeled batch size $B$, the number of categories $C$, and the visual dimension $H\times W$, the loss weight $\omega$ for each class can be formulated though eq. 1:
$\omega_{c}=1-\sqrt{\frac{\sum_{b=1}^{B}\sum_{h=1}^{H}\sum_{w=1}^{W}I({M}_{b,c,h,w}=1)}{B\times H\times W}}$ (1)
The $I(\cdot)$ represents the indicator function, which returns 1 as the condition is true and 0 otherwise. Specifically, $I({M}_{b,c,h,w}=1)$ indicates whether a pixel located at position $(h,w)$ in the $b$-th image of a batch belongs to category $c$. The calculated category weights are then normalized to ensure the sum of the weights equals 1 using eq. 2:
$\omega_{c}^{\prime}=\frac{\omega_{c}}{\sum_{c=1}^{C}\omega_{c}}$ (2)
@@
$\lambda(t,t_{max})=\begin{cases}1&t_{max}=0\\ \exp\left(-5\left(1-\frac{t}{t_{max}}\right)^{2}\right)&0\leq t\leq t_{max}\end{cases}$ (3)
@@ 4.1 Dataset Preprocessing
We resize and crop the images and masks to 224 $\times$ 224 resolution. As the tube location varies and holds a small size, directly leveraging the center crop may lead to the entire tube being cropped out. To this end, we propose a dynamic crop approach based on the tube location. Firstly, we resize the images and mask and keep the aspect ratio unchanged to ensure the minimum side equals 224. Secondly, we detect the tube location and set the average tube position to the crop center. The center of the image and mask are regarded as the crop center in case the tube is not detected. Finally, we adjust the location of the crop area to ensure that the cropped area is not out of bounds and perform the crop.
@@ 4.2 Configuration and Evaluation
We conduct extensive segmentation and classification experiments to provide comprehensive and convincing evaluations. For segmentation, we segment the four structures in the UNGT dataset, including the liver, stomach, tube, and pancreas. Regarding classification, we classify the existence of the tube within each image. We implement our experiments with NVIDIA RTX A6000 and Intel Xeon 3204 and leverage PyTorch as the training framework. We report the mean and standard deviation of three independent runs with varying seeds in percentage and present the standard deviation in superscript.
Regarding segmentation, we one-hot encode the mask, ensuring the tube location is encoded to both the stomach and tube channel. We partite the dataset into a training and a validation subset with a ratio of 7:3 and leverage 80% labeled and 20% unlabeled images to train semi-supervised approaches. We set SGD as the optimizer and assign the momentum and weight decay to 0.9 and 0.0001, respectively. We assign the initial learning rate to 0.01. We leverage the cosine annealing schedule [46] to update the learning rate during training AAMS. We set the maximum training iterations to 20000 and the batch size to 8, with an identical labeled and unlabeled batch size of 4 for semi-supervised approaches. The loss functions and data augmentation strategies vary across methods. We evaluate the model performance leveraging dice score (DSC), sensitivity (SEN), and precision (PRE).
Considering classification, we divide the dataset into a training subset and a validation subset with a ratio of 7:3. We maintain the ratio of examples with and without the tube in the training and validation subsets, ensuring the class balance across each category. We leverage Adam as the optimizer and set the weight decay to 0.001. We assign the initial learning rate to 0.01 and decrease it with a factor of 0.5 and patience of 5 as training proceeds. We set the maximum epoch to 200 and the batch size to 16. We leverage the cross-entropy loss as the loss function and perform data augmentation through the OAA. Metrics for performance evaluation include accuracy (ACC), SEN, PRE, specificity (SPE), negative predictive value (NPV), and F-1 score (FOS).
@@ 5.1 Segmentation
Table 1: Segmentation results across AAMS and SOTA approaches on the developed UNGT dataset. L, S, T, and P present the liver, stomach, tube, and pancreas, respectively.
@@
We present the segmentation results across AAMS and SOTA methods on the developed UNGT dataset in table 1. As can be seen, the introduced AAMS represents superior performance leadership compared with SOTA methods to varying extents on four structures. Regarding the liver, the AAMS achieves the highest DSC and PRE of 72.74% and 77.94%. Though CTCT presents the optimal SEN of 73.79%, it does not exhibit an extraordinary PRE of 74.84%. The MedT does not present exemplary performance under the current configuration with a DSC of 59.85%. Considering the stomach, the AAMS realizes the optimal DSC and SEN of 78.55% and 81.67%. Although MGCC archives the highest PRE, it achieves a relatively low SEN of 75.39%. Analogously, the MedT demonstrates an unexpected performance with a DSC equal to 63.91%. Significant performance decreases are identified in the results of the tube segmentation, preeminently attributed to the unbalanced constraint within the dataset. Despite this, the AAMS realizes lower performance degradation across varying approaches and achieves the optimal DSC, SEN, and PRE of 41.07%, 42.63%, and 43.04%. Greater performance degradation is observed for most SOTA methods and MedT realizes an extremely low DSC of merely 4.16%. Analogously, further performance decreases are observed in pancreas segmentation due to the severe limitation on the number of images. Despite this, the AAMS realizes the optimal DSC and PRE of 20.74% and 23.19%. Though CCT achieves the optimal SEN of 20.04%, AAMS follows closely with a performance of 19.98%. The lowest DSC of 9.44% is achieved by MedT, demonstrating higher performance compared with the tube. This suggests that MedT may not be well-suited for segmenting extremely minor objects. The performance difference among varying structures reflects the difficulty of segmenting the unbalanced UNGT dataset and the metrics leadership of the proposed AAMS emphasizes its adaptability and variability.
@@
Figure 3: Predicted and GT masks across AAMS and SOTA approaches on the introduced UNGT dataset. GT stands ground truth.
@@
We demonstrate the predicted and GT masks across AAMS and SOTA approaches on the introduced UNGT dataset in fig. 3. Results indicate that the AAMS demonstrates boosted prediction accuracy compared with varying SOTA approaches. In the first image, the AAMS effectively prevents the merging of adjacent structural areas and ensures accurate delineation, while many approaches such as ICT output the liver and stomach areas without a clear boundary. For the fifth image, the AAMS predicts the most comprehensive pancreas areas while several approaches such as CCT and R-Drop present a fragmented pancreas. Concerning the seventh image, the AAMS outputs the most accurate stomach shape though the stomach shape deviates from common forms. Meantime, several approaches such as UNext and URPC predict undesirable stomach shapes. As for the eighth figure, the AAMS predicts the tube and pancreas areas more accurately, whereas approaches such as UNext yield an imprecise tube, and CMUNeXt outputs a fragmented pancreas. Analogously, the MedT demonstrates the most deficient prediction areas. A marked proportion of its prediction regions are fragmented, featuring unreasonable structure boundaries identified. For instance, rectangle boundaries are observed within the stomach structure in the seventh and the last images. It is worth noting that most of the approaches tend to wrongly predict the light region as a tube due to their analogous characteristics. An intuitive example is shown in the last image, where the light-long strip region is erroneously identified as a tube by most methods. The vast number of wrongly predicted regions emphasizes the difficulty of segmenting the introduced UNGT dataset and the relatively superior prediction accuracy demonstrates the effectiveness of the developed AAMS.
@@
We visualize the cumulative distribution function across different structures regarding DSC for varying methods on the developed UNGT dataset in fig. 4. The red dotted lines represent the proportion of the ground truth (GT) mask in which the corresponding structure does not exist in percentage. The predictions below the red dotted lines may not be explained as inaccurate as the DSC can equal zero in case both GT and prediction masks do not present the structure. Through identification, it can be found that the AAMS demonstrates superior overall performance across the four structures. Considering the liver, AAMS presents as one of the optimal approaches, while MedT and UNext reveal relatively undesirable results. Similar observations can be seen in the results of stomach segmentation, where most methods demonstrate desirable performance except for MedT and UNext. Concerning the tube, the AAMS demonstrates tremendous performance leadership compared with the remaining approaches. Under this challenging scenario, the performance of MedT and UNext deteriorates dramatically and most of the tube prediction of MedT presents a DSC of zero. The trends of the pancreas conform to that of the tube, in which the AAMS demonstrates apparent leadership while MedT and UNext fall behind. The undesirable results presented in the tube and pancreas segmentation reveal the difficulty in segmenting the unbalanced UNGT dataset, and the substantial performance leadership on these structures illustrates the priority of the introduced AAMS.
@@
We present the Bland-Altman plot of AAMS across varying structures on the introduced UNGT dataset in fig. 5. We calculate the ratio of each structure pixel in the predicted and GT masks by counting the number of pixels and dividing it by $224^{2}$. The consistency across each prediction-GT pair is then compared. From the results, the prediction of AAMS depicts superior consistency compared with the GT. The mean differences are overall close to 0%, with the maximal and minimal differences equal to 0.33% and 0.05%, for the liver and stomach respectively. The standard deviations lie within the expected range and vary with the size of different structures. Explicitly, the highest and lowest standard deviations of 5.85% and 0.40% are identified for the stomach and tube. Besides, a dominant portion of points is located within the 95% limits of agreement. Though several outliers are observed for each structure, the number of outliers aligns with the expectation and can be attributed to the limited size of the UNGT dataset. In a dataset with a few hundred images, certain patterns may be present in the validation subset but absent or underrepresented in the training subset. Notably, most of the outliers of the tube and pancreas lie along the negative axis, indicating that the AAMS consistently tends to underpredict these structures. This observation conforms to our expectations since most of these structures are relatively small, leading the network to underestimate the sizes of larger tubes and pancreas. Under this case, the difference across predicted and GT masks is primarily influenced by the structure-specific attributes instead of the inherent constraints of the AAMS. Besides, plots across varying structures evince no uniform difference trend, demonstrating the observed discrepancies lack a stringent systematic pattern. The high consistency among the prediction-GT mask pairs reveals the effectiveness of the proposed AAMS.

### v2

@@ 3.1 Data Acquisition and Characteristics
To mitigate the gap from the dataset aspect, we propose a UNGT dataset with 493 US images gathered from 110 patients. We demonstrate several example images within the developed UNGT dataset in fig. 1a. For neatness considerations, images have been resized and cropped as discussed in Section 4.1. Of these, 305 images are captured in the transverse section, and 188 are scanned in the longitudinal section. The average resolution of the scans is approximately 879 $\times$ 583. The UNGT dataset comprises precise annotations for the liver, stomach, tube, and pancreas, and we partition it into two components. Specifically, component A encompasses 321 images with the tube collected from 60 patients, and component B consists of 172 images without the tube gathered from 50 patients.
@@
We construct the dataset in two ways. First, we retrospectively trawl the institutional PACS acquired between 1 January 2020 and 31 December 2022 to retrieve pre-existing US images demonstrating the stomach with or without a feeding tube. Second, we prospectively recruit patients with a pre-existing feeding tube in situ for US scanning after the approval of the ethics review board and the acquisition of written patient consent. The IRB number for our work is 2022/00171. We perform the scans using standard commercial US machines, including General Electric LOGIQ E9 and E10, Siemens ACUSON, and Toshiba APLIO 100 on transverse and longitudinal sections. We adjust the curvilinear probe frequency between 2–8 MHz based on the patient’s habitus, gain, depth, and focal point on the image for optimal image contrast and resolution. Incorporating images from different US machines introduces heterogeneity into the dataset, enhancing its robustness and real-world applicability. We obtain approximately 2 to 10 static images per patient and capture more frames when the cine-clips are applicable. We review the quality of the US images to ensure adequate clarity, resolution, and field-of-view before selection. A board-certified radiologist with 10 years of experience in US imaging selects the scans, performs cropping to remove confidential information, and carries out the annotation using an open-source tool Labelme [45]. The liver, stomach, tube, and pancreas are annotated, with the tube consistently inside the stomach. Another board-certified radiologist with 5 years of experience in US imaging cross-checks the annotations for accuracy. Annotation discrepancies are resolved by consensus between the two radiologists.
@@
We present the characteristics analysis of our developed UNGT dataset in fig. 1b to fig. 1e. The distribution of image counts per patient reveals that only a limited number of images are typically captured per patient, highlighting the challenges of US image acquisition. The brightness histogram distribution aligns with the typical characteristics of US images, which often exhibit predominantly low average brightness. We exclude the grey level of zero as it can comprise non-scanned regions, such as black triangle borders outside the scan boundary. The primary challenge within the UNGT dataset is the severely unbalanced distribution. Considering the image counts per structure, most images comprise the liver and stomach, while only a small portion involves the tube and pancreas. Such an unbalanced constraint becomes more severe for the size across varying structures. The number of pixels labeled as liver and stomach is not at the same magnitude as that of the tube and pancreas. This can be attributed to the limited number of structures and the small size of such objects. The severely unbalanced distribution at coarse and fine levels demonstrates the difficulty in analyzing the UNGT dataset, especially for the tube and pancreas structures.
@@ 3.2 Model Architecture Design
Figure 2: Architecture of the developed AAMS. It comprises an ADW scheme, an MAA block, the OAA, and the RRE blocks. The loss function of AAMS includes fusion loss $L_{f}$, sensitivity loss $L_{s}$, and unsupervised loss $L_{u}$. The ADW scheme calculates each class’s normalized loss weight $\omega^{\prime}$. The MAA block is mainly composed of a local excitation (LEX) block, a global excitation (GEX) block, and a multi-layer perceptron (MLP) block. $\oplus$ and $\times$ denote the exclusive OR and multiplication operations. L, S, T, and P stand for the liver, stomach, tube, and pancreas. HAAP and VAAP represent horizontal and vertical adaptive average pooling layers. DwConv, SpConv, and PwConv present depthwise, stripe, and pointwise convolution layers. BN stands for batch normalization layer. (a) AAMS architecture, (b) ADW scheme, and (c) MAA block.
@@
We advance the DEMS to the multi-category segmentation context using parallel binary segmentation and propose an ADW scheme and an MAA block. We present the detailed architecture of the proposed AAMS in fig. 2. It comprises an ADW scheme, an MAA block, the OAA, and the RRE blocks. The loss function of AAMS comprises fusion loss $L_{f}$, $L_{s}$, and unsupervised loss $L_{u}$. The $L_{f}$ is composed of BCE loss and dice loss, while the $L_{u}$ is based on MSE loss. The ADW scheme tackles class imbalance by adaptively regulating the loss contributions across various categories during training. The MAA block enhances feature representation by integrating local and global contextual representations.
The ADW scheme computes the number of pixels for each category for each batch. Given a batch of encoded mask $M$, the labeled batch size $B$, the number of categories $C$, and the visual dimension $H\times W$, the loss weight $\omega$ for each class can be formulated though eq. 1:
$\omega_{c}=1-\sqrt{\frac{\sum_{b=1}^{B}\sum_{h=1}^{H}\sum_{w=1}^{W}I({M}_{b,c,h,w}=1)}{B\times H\times W}}\,$, (1)
where $I(\cdot)$ represents the indicator function, which returns 1 if the condition is true and 0 otherwise. Specifically, $I({M}_{b,c,h,w}=1)$ indicates whether a pixel located at position $(h,w)$ in the $b$-th image of a batch belongs to category $c$. The calculated category weights are then normalized to ensure the sum of the weights equals 1 using eq. 2:
$\omega_{c}^{\prime}=\frac{\omega_{c}}{\sum_{c=1}^{C}\omega_{c}}\,.$ (2)
@@
$\lambda(t,t_{max})=\begin{cases}1&t_{max}=0\,,\\ \exp\left(-5\left(1-\frac{t}{t_{max}}\right)^{2}\right)&0\leq t\leq t_{max}\,.\end{cases}$ (3)
@@ 4.1 Dataset Preprocessing
We resize and crop the images and masks to 224 $\times$ 224 resolution. As the tube location varies and holds a small size, directly leveraging the center crop may lead to the entire tube being cropped out. To this end, we propose a dynamic crop approach based on the tube location. Firstly, we resize the images and mask and keep the aspect ratio unchanged to ensure the minimum side equals 224. Secondly, we detect the tube location and set the average tube position to the crop center. The center of the image and mask is regarded as the crop center in case the tube is not detected. Finally, we adjust the location of the crop area to ensure that the cropped area is not out of bounds and perform the crop.
@@ 4.2 Configuration and Evaluation
We conduct extensive segmentation and classification experiments to provide comprehensive and convincing evaluations. For segmentation, we segment the four structures in the UNGT dataset, including the liver, stomach, tube, and pancreas. Regarding classification, we classify the existence of the tube within each image. We leverage PyTorch as the training framework. We report the mean and standard deviation of three independent runs with varying seeds in percentage and present the standard deviation in superscript.
Regarding segmentation, we one-hot encode the mask, ensuring the tube location is encoded to both the stomach and tube channel. We partite the dataset into a training and a validation subset with a ratio of 7:3 and leverage 80% labeled and 20% unlabeled images to train semi-supervised approaches. We set SGD as the optimizer and assign the momentum and weight decay to 0.9 and 0.0001. We assign the initial learning rate to 0.01. We leverage the cosine annealing schedule [31] to update the learning rate during training AAMS. We set the maximum training iterations to 20000 and the batch size to 8, with an identical labeled and unlabeled batch size of 4 for semi-supervised approaches. The loss functions and data augmentation strategies vary across methods. We evaluate the model performance leveraging dice score (DSC), sensitivity (SEN), and precision (PRE).
Considering classification, we divide the dataset into a training subset and a validation subset with a ratio of 7:3. We maintain the ratio of examples with and without the tube in the training and validation subsets, ensuring the class balance across each category. We leverage Adam as the optimizer and set the weight decay to 0.001. We assign the initial learning rate to 0.01 and decrease it by a factor of 0.5 and patience of 5 as training proceeds. We set the maximum epoch to 200 and the batch size to 16. We leverage the cross-entropy loss as the loss function and perform data augmentation through the OAA. Metrics for performance evaluation include accuracy (ACC), SEN, PRE, specificity (SPE), negative predictive value (NPV), and F-1 score (FOS).
@@ 5.1 Segmentation
Table 1: Segmentation results across AAMS and SOTA approaches on the developed UNGT dataset. L, S, T, and P present the liver, stomach, tube, and pancreas. denotes semi-supervised approach.
@@
We present the segmentation results across AAMS and SOTA methods on the developed UNGT dataset in table 1. As can be seen, the introduced AAMS represents superior performance leadership compared with SOTA methods to varying extents on four structures. Regarding the liver, the AAMS achieves the highest DSC and PRE of 72.74% and 77.94%. Though CTCT presents the optimal SEN of 73.79%, it does not exhibit an extraordinary PRE of 74.84%. The MedT does not present exemplary performance under the current configuration with a DSC of 59.85%. Considering the stomach, the AAMS realizes the optimal DSC and SEN of 78.55% and 81.67%. Although MGCC archives the highest PRE, it achieves a relatively low SEN of 75.39%. Analogously, the MedT demonstrates an unexpected performance with a DSC equal to 63.91%. Significant performance decreases are identified in the results of the tube segmentation, preeminently attributed to the unbalanced constraint within the dataset. Despite this, the AAMS realizes lower performance degradation across varying approaches and achieves the optimal DSC, SEN, and PRE of 41.07%, 42.63%, and 43.04%. Greater performance degradation is observed for most SOTA methods, and MedT realizes an extremely low DSC of merely 4.16%. Analogously, further performance decreases are observed in pancreas segmentation due to the severe limitation on the number of images. Despite this, the AAMS realizes the optimal DSC and PRE of 20.74% and 23.19%. Though CCT achieves the optimal SEN of 20.04%, AAMS follows closely with a performance of 19.98%. The lowest DSC of 9.44% is achieved by MedT, demonstrating higher performance compared with the tube. This suggests that MedT may not be well-suited for segmenting extremely minor objects. The performance difference among varying structures reflects the difficulty of segmenting the unbalanced UNGT dataset, and the metrics leadership of the proposed AAMS emphasizes its adaptability and variability.
@@
Figure 3: Predicted and GT masks across AAMS and SOTA approaches on the introduced UNGT dataset. GT stands for ground truth.
@@
We demonstrate the predicted and ground truth (GT) masks across AAMS and SOTA approaches on the introduced UNGT dataset in fig. 3. Results indicate that the AAMS demonstrates boosted prediction accuracy compared with varying SOTA approaches. In the first image, the AAMS effectively prevents the merging of adjacent structural areas and ensures accurate delineation, while approaches like ICT output the liver and stomach areas without a clear boundary. For the fifth image, the AAMS predicts the most comprehensive pancreas areas, while approaches such as CCT and R-Drop present a fragmented pancreas. Concerning the seventh image, the AAMS outputs the most accurate stomach shape despite the stomach shape deviating from common forms. Meanwhile, models such as UNext and URPC predict undesirable stomach shapes. As for the eighth figure, the AAMS predicts the tube and pancreas areas more accurately, whereas approaches such as UNext yield an imprecise tube, and CMUNeXt outputs a fragmented pancreas. Analogously, the MedT demonstrates the most deficient prediction areas. A marked proportion of its prediction regions are fragmented, featuring unreasonable structure boundaries identified. For instance, rectangle boundaries are observed within the stomach structure in the seventh and last images. It is worth noting that most of the approaches tend to wrongly predict the light region as a tube due to their analogous characteristics. An intuitive example is shown in the last image, where the light-long strip region is erroneously identified as a tube by most methods. The vast number of wrongly predicted regions emphasizes the difficulty of segmenting the introduced UNGT dataset, and the relatively superior prediction accuracy demonstrates the effectiveness of the developed AAMS.
@@
We visualize the cumulative distribution function across different structures regarding DSC for varying methods on the developed UNGT dataset in fig. 4. The red dotted lines represent the proportion of the GT mask in which the corresponding structure does not exist in percentage. The predictions below the red dotted lines may not be explained as inaccurate as the DSC can equal zero in case both GT and prediction masks do not present the structure. Through identification, it can be found that the AAMS demonstrates superior overall performance across the four structures. Considering the liver, AAMS presents as one of the optimal approaches, while MedT and UNext reveal relatively undesirable results. Similar observations can be seen in the results of stomach segmentation, where most methods demonstrate desirable performance except for MedT and UNext. Concerning the tube, the AAMS demonstrates tremendous performance leadership compared with the remaining approaches. Under this challenging scenario, the performance of MedT and UNext deteriorates dramatically, and most of the tube prediction of MedT presents a DSC of zero. The trends of the pancreas conform to that of the tube, in which the AAMS demonstrates apparent leadership while MedT and UNext fall behind. The undesirable results presented in the tube and pancreas segmentation reveal the difficulty in segmenting the unbalanced UNGT dataset, and the substantial performance leadership on these structures illustrates the priority of the introduced AAMS.
@@
We present the Bland-Altman plot of AAMS across varying structures on the introduced UNGT dataset in fig. 5. We calculate the ratio of each structure pixel in the predicted and GT masks by counting the number of pixels and dividing it by $224^{2}$. The consistency across each prediction-GT pair is then compared. From the results, the prediction of AAMS depicts superior consistency compared with the GT. The mean differences are overall close to 0%, with the maximal and minimal differences equal to 0.33% and 0.05% for liver and stomach. The standard deviations lie within the expected range and vary with the size of different structures. Explicitly, the highest and lowest standard deviations of 5.85% and 0.40% are identified for the stomach and tube. Besides, a dominant portion of the points is located within the 95% limits of agreement. Though several outliers are observed for each structure, the number of outliers aligns with the expectation and can be attributed to the limited size of the UNGT dataset. In a dataset with a few hundred images, certain patterns may be present in the validation subset but absent or underrepresented in the training subset. Notably, most of the outliers of the tube and pancreas lie along the negative axis, indicating that the AAMS consistently tends to underpredict these structures. This observation conforms to our expectations since most of these structures are relatively small, leading the network to underestimate the larger ones. In this case, the difference across predicted and GT masks is primarily influenced by the structure-specific attributes instead of the inherent constraints of the AAMS. Besides, plots across varying structures evince no uniform difference trend, demonstrating that the observed discrepancies lack a stringent systematic pattern. The high consistency among the prediction-GT mask pairs reveals the effectiveness of the proposed AAMS.